OpenStudy (anonymous):

Find the inverse of the function. f(x)=(3/-x-3)+1 (the 3 is actually over the -x-3)

OpenStudy (anonymous):

\[f(x)=\frac{3}{-x-3}+1\]?

OpenStudy (anonymous):

you can start by writing \[x=\frac{3}{-y-3}+1\] and solve for \(y\)

OpenStudy (anonymous):

any idea how to do that?

OpenStudy (anonymous):

No, that first step is what I thought I needed to do but I haven't touched this material in 8 months.

OpenStudy (anonymous):

you can do a few different things a good first step might be to subtract 1 from both sides, get \[x-1=\frac{3}{-y-3}\] then flip and and get \[\frac{3}{x-1}=-y-3\]

OpenStudy (anonymous):

add 3 and then change the sign to finish

OpenStudy (anonymous):

oh, ok. Thanks so much =)